Please prove it


1/a + 1/b +1/c + 1/d >= 64/(a+b+c+d)

Grade:12

1 Answers

jitender lakhanpal
10 years ago

apply A.M>=G.M   for a,b,c,d    and for 1/a,1/b,1/c,1/d

we get     1/(abcd)^(1/4) > = 4/(a+b+c+d)

{(1/a) + (1/b) +(1/c) + (1/d)}/4 >=1/{(abcd)^4}

combining we get

the given equation

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)